I need to write in simplest form
3 5/12 - 1 7/12 =
THanks

Found 4 solutions by stanbon, trazana, Fombitz, MusicalK10:
Answer by stanbon(75887)

3 5/12 - 1 7/12 =
= 41/12 - 19/12
= 22/12
= 11/6
= 1 5/6

===========
Cheers,
Stan H.

Answer by trazana(1)
Answer by Fombitz(32388)

Change your mixed numbers to fractions.
3 5/12=41/12
1 7/12=19/12
3 5/12 - 1 7/12 = 41/12-19/12
Same denominator, you can subtract.
3 5/12 - 1 7/12 = 41/12-19/12
3 5/12 - 1 7/12 = (41-19)/12
3 5/12 - 1 7/12 = 22/12
3 5/12 - 1 7/12 = (2*11)/(2*6)
3 5/12 - 1 7/12 = 11/6
3 5/12 - 1 7/12 = 1 5/6

Answer by MusicalK10(1)

3 5/12 - 1 7/12 =
3 5/12 = 3 x 12 + 5 = 41 or 41/12
1 7/12 = 1 x 12 + 7 = 19 or 19/12
41/12 - 19/12 = 22/12
22/12 divided by 2/2 = 11/6
11 divided by 6 = 1 R5 which equals 1 5/6

Answer = 1 5/6
Hope this helps.
-MusicalK10